hello,
i am having great difficulty with trigonometry. can u please help by providing a step by step solution to the following question. thank you very much.
Question:
A ship sails 12 km due north of a port and then sails 14 km due east.
(a) How far is the ship from the port?
(b) How much further east will it have sailed when it is 30 km from the port?

Answer by richwmiller(17219)

Actually no trig involved here.
Pure geometry
a)Right triangle as a matter of fact. 12 and 14 are the legs
12^2+14^2=c^2
Find the hypotenuse.
b)New triangle with hypotenuse 30 and one leg 12
12^2+x^2=30^2
Find the leg going east
discard negative answer
x-14 will be the answer
27.49-14=13.49 km to go
or do it this way
12^2+(14+y)^2=30^2
y will be the answer.
discard the negative answer
13.49 km more